(i) A singly generated Banach algebra (ii) A Banach algebra can not be singly generated

******************************

(i) C_{∘}((0, 1]) is singly generated by the inclusion function t ⟼ t, by the Stone-Weierstrass theorem.

(ii) C(Γ), where Γ is the unit circle in plane.